# Bilinear relational structure fixes reversal curse and enables consistent model editing

### v2

@@ Abstract
The reversal curse—a language model’s (LM) inability to infer an unseen fact “B is A” from a learned fact “A is B”—is widely considered a fundamental limitation. We show that this is not an inherent failure but an artifact of how models encode knowledge. By training LMs from scratch on a synthetic dataset of relational knowledge graphs, we demonstrate that bilinear relational structure emerges in their hidden representations. This structure substantially alleviates the reversal curse, enabling LMs to infer unseen reverse facts. Crucially, we also find that this bilinear structure plays a key role in consistent model editing. When a fact is updated in a LM with this structure, the edit correctly propagates to its reverse and other logically dependent facts. In contrast, models lacking this representation not only suffer from the reversal curse but also fail to generalize edits, further introducing logical inconsistencies. Our results establish that training on a relational knowledge dataset induces the emergence of bilinear internal representations, which in turn enable LMs to behave in a logically consistent manner after editing. This implies that the success of model editing depends critically not just on editing algorithms but on the underlying representational geometry of the knowledge being modified.
@@ 1 Introduction
Language models (LMs) have become powerful tools for knowledge-intensive tasks, yet their reasoning capabilities often fall short of human-level logical consistency (Berglund et al., 2024; Allen-Zhu & Li, 2025); a prominent example is the reversal curse: a model trained on “A is the parent of B” frequently fails to infer the reverse fact, “B is the child of A.” This failure suggests that LMs learn shallow, directional associations rather than robust, symmetrical relationships, undermining their reliability. Ensuring logical consistency is particularly challenging in model editing, which seeks to update factual knowledge in a trained model without costly retraining from scratch. (De Cao et al., 2021; Meng et al., 2022). An ideal edit should propagate logically; for instance, changing “A is the spouse of C” to “A is the spouse of D” should automatically imply that “D is the spouse of A” and “B is the child of D.”
However, existing approaches struggle with this logical generalization. Model editing methods often fail to propagate updates to the entailed facts, requiring that both directions of a relationship be explicitly co-edited to avoid the reversal curse (Thibodeau, 2022; Yao et al., 2023; Hase et al., 2024). This limitation raises a critical question: Are these failures an inherent flaw in the transformer architecture, or are they an artifact of how models learn to represent knowledge? While prior efforts have focused on explaining the directional nature of autoregressive objectives (Zhu et al., 2024; Kitouni et al., 2024), the latter question—whether the geometry of internal knowledge representations is responsible—remains unexplored.
This work begins with the hypothesis that logical failures in a LM may stem from the model’s internal knowledge representation. Accordingly, we investigated how relational knowledge is encoded using three knowledge representation probes: linear, translational, and bilinear probes (see Figure 1). Recent work has demonstrated that knowledge decoding from a pretrained LM’s internal representation can be approximated by either linear mapping (Hernandez et al., 2024) or translational mapping (Merullo et al., 2024). While these frameworks offer valuable insights, they fall short in capturing symmetric and compositional relations, which are essential for robust logical reasoning. We instead focus on bilinear relational models such as RESCAL (Nickel et al., 2011) in knowledge graph embedding literature, which represent relations as matrices that mediate interactions between entities. Bilinear relational structures naturally capture inverse relations (via matrix transposition) and compose relations (via matrix multiplication), providing a rich algebraic framework for reasoning.
@@
  • We demonstrate that with appropriate regularization (weight decay), LMs can overcome the reversal curse by learning a robust bilinear relational structure, achieving near-perfect accuracy on unseen reverse relations (Figure 2, right).
  • Using multiple probes on LM hidden representations, we find that a bilinear probe best explains them, with the signal emerging in intermediate layers (Figure 3). The learned relation matrices also satisfy composition and inversion tests (Figure 4), supporting the presence of a bilinear structure in LMs.
  • We find a strong association between this bilinear structure and editing generalization. Models possessing this structure successfully propagate edits to logically related facts, whereas models lacking it fail to generalize, despite the fact that the direct edit is successful (Figure 5).
Building on these findings, our study introduces a substantial change in perspective: the key to resolving logical failures lies not solely in the model architecture or editing algorithm, but fundamentally on the geometric structure of learned knowledge representations. This result suggests the pivotal role of relational encoding in shaping the reliability and robustness of model behavior.
Figure 1: Schematics of the three relational embedding structures examined in our study. Given a subject $s$ and an object $o$, the relation $r$ can be represented as (a) a linear transformation, (b) a vector translation, or (c) a bilinear interaction mediated by a relation-specific matrix $M_{r}$. In a fact “son of Tracy Lance Smith is Cory Lance Smith,” the subject $s$ is “Tracy Lance Smith,” the object $o$ is “Cory Lance Smith,” and the relation $r$ is son.
@@ 2 Related work
The reversal curse—the failure to infer “B is A” from “A is B”—has been identified as a fundamental limitation of LMs (Berglund et al., 2024). Prevailing hypotheses attribute this to the directional nature of the autoregressive training objective, which preferentially models $P(\text{B}|\text{A})$ but not $P(\text{A}|\text{B})$ (Allen-Zhu & Li, 2025; Zhu et al., 2024; Kitouni et al., 2024). The nature of this failure is nuanced; Lin et al. (2024) suggest that it may be an issue of knowledge retrieval rather than storage, as models that fail open-ended generation can succeed on multiple-choice questions where the answer is present in a prompt. Proposed solutions often directly target the training process. These include data augmentation via subword-level reordering in a sentence (Golovneva et al., 2024), generating reversed examples by LMs (Lampinen et al., 2025), or modifying the training objective to be direction-agnostic (Kitouni et al., 2024). While these studies offer practical mitigation, they often treat it as an unavoidable artifact of the training objective. Our work offers a different perspective by focusing on the representations in LMs.
@@
Model editing aims to update facts in LMs without costly retraining (De Cao et al., 2021). Various model editing methods have been proposed, such as ROME (Meng et al., 2022) and MEMIT (Meng et al., 2023), which show better generalization than naive fine-tuning, but all editing algorithms suffer from the reversal curse (Thibodeau, 2022; Yao et al., 2023). This points to a deeper conceptual challenge. Hase et al. (2024) provide a comprehensive critique, framing model editing as an instance of belief revision, for which no simple solution exists. They argue that logical generalization is not just a desirable feature but a core requirement for any successful editing paradigm, and they demonstrate empirically that current methods often fail to produce coherent belief updates. While these works call for better testbeds, the underlying mechanisms causing these failures remain an open question. In our work, instead of proposing new editing algorithms, we investigate the prerequisite internal structures that enable logical consistency after model editing.
Mathematical Structures for Relational Knowledge in LMs.
A substantial body of research has investigated where LMs store factual knowledge (Geva et al., 2021; Meng et al., 2022; Pan et al., 2025). The mechanism for retrieving specific facts is complex and distributed across multiple layers and attention heads, as revealed by studies of interventions (Hase et al., 2023) and attention mechanisms (Geva et al., 2023). In contrast, far fewer studies examine which mathematical structure LMs employ to resolve relation. Recent investigations suggest that the underlying structures are unexpectedly simple: Hernandez et al. (2024) demonstrate that LMs implicitly implement Linear Relational Embeddings (Paccanaro & Hinton, 2002), while Merullo et al. (2024) show that they exploit translational structures familiar from Word2Vec (Mikolov et al., 2013). However, the attention mechanism in LMs is fundamentally built on more expressive bilinear operations (Elhage et al., 2021). Furthermore, such bilinear models have a long history of success in modeling relational data, particularly in knowledge graph embedding methods like RESCAL (Nickel et al., 2011). Despite bilinear operations being central to transformer architecture and successful in relational learning, no work has investigated whether LMs exploit bilinear structures for decoding relational knowledge. This paper presents the first systematic analysis of the bilinear structures underlying relational knowledge in LMs.
@@ 3 Relational Knowledge Dataset and Language Models
In order to test our hypothesis, we create a synthetic relational knowledge dataset and train multiple LMs with different hyperparameters from scratch on it.
Figure 2: (Left) A schematic of the synthetic family knowledge graph used for experiments. Nodes represent entities, and edges represent one of eight relations. (Right) Test accuracy on the unseen relations (mother/father) as a function of weight decay. Each weight decay setting was trained using three different random seeds.
Synthetic Knowledge Graph and Task
@@
Language models
We train decoder-only transformers using the GPT-NeoX architecture (Andonian et al., 2023) from scratch on the synthetic dataset. Each model has 12 layers, a hidden size of 896, and 16 attention heads, totaling approximately 206M parameters. We employ this architecture in the following sections. More details of architectural and training are provided in the Appendix A.
@@ 4 Experiments
Our experiments seek to uncover the internal mechanisms that enable LMs to perform logical reasoning. We first demonstrate that even when the training data contain sufficient information to infer reverse relations, models only overcome the reversal curse when guided by appropriate regularization (Experiment 1). We then pivot to the central question of our work: which mathematical structure enables this success? Through a series of probing experiments, we uncover an emergent bilinear relational structure in the successful models (Experiment 2) and verify its algebraic properties (Experiment 3). Finally, we demonstrate that this geometry plays a central role in addressing model editing challenges. We show that the bilinear structure is a key to ensuring that edits propagate in a logically consistent manner, establishing a unified explanation for both the reversal curse and editing generalization failures (Experiment 4).
4.1 Experiment 1: Training LMs on relational knowledge dataset
We test whether a LM can learn to infer withheld relations (father/mother) by observing their reverse counterparts (e.g., son/daughter) and compositional examples elsewhere in the training set. We use AdamW (Loshchilov & Hutter, 2019) with a learning rate of $3\times 10^{-4}$ and sweep weight decay over $\{0,0.1,0.5,1.0,2.0,3.0,4.0,5.0,6.0\}$, training three random seeds per setting (27 models in total).
The results, shown in Figure 2 (right), reveal a striking dependency. Without sufficient regularization, models consistently fail, showing the reversal curse with low accuracy (weight decay<~<~1.0). However, as weight decay is increased, a split in outcomes emerges: some models remain “reversal cursed,” while others break the curse and achieve near-perfect accuracy on the unseen reverse facts. This transition indicates that the reversal curse is not an inherent limitation but an artifact of an under-constrained training objective, where regularization promotes a more generalizable internal structure over simple memorization.
@@ 4.2 Experiment 2: Probing internal representation for relational structures
This probe tests for a linear relational structure (Paccanaro & Hinton, 2002) in LMs. Hernandez et al. (2024) models a relation $r$ as the local affine transformation that the transformer applies to map a subject representation $s_{l}$ from layer $l$ to the object’s pre-prediction representation $o_{L}$ in the final layer $L$. This yields the approximation $o_{L}\approx W_{r}s_{l}+b_{r}$, where $o_{L}$ is the hidden state at the position immediately preceding the object token (see Figure 1a).
@@
Due to the high computational cost of calculating the Jacobians, we estimate these parameters by averaging a small sample of $n=10$ training examples per relation.
@@
This probe, which tests for the kind of vector arithmetic investigated in LMs by Merullo et al. (2024), models a relation $r$ as a simple vector offset. For each relation $r$, we fit a translation vector $v_{r}$ such that $s_{l}+v_{r}\approx o_{l}$. Unlike the linear relation embedding, $o_{l}$ is taken from the same layer $l$ as $s_{l}$ (see Figure 1b) at the last token position of the object name.
@@
between any entity given the relation $r$ where matrix $M_{r}\in\mathbb{R}^{d\times d}$. The target function is:
@@ Results.
For each layer $l$ and each relation $r$, we trained a separate probe and evaluated its accuracy on testset entities (Figure 3). The gray dashed line at 1/3 marks a chance-level baseline. In our family graph (see Figure 2, left), for any relation $r$ there are only three candidates per family that satisfy $f^{*}_{r}(s,o)=1$. A probe that can recognize the family name and the relation $r$ but fails to use the subject embedding would therefore guess uniformly among these three candidates, yielding an expected accuracy of $1/3$. For example, given the text “[Subject] Scott Wall husband [Object]”, a probe that ignores the subject can narrow the object to the three husband candidates in the “Scott Wall” family and would be correct one out of three times on average.
@@
Figure 3: Layer-wise averaged accuracy of relational embedding probes across all relations for “Reversal Cursed” models (blue) and “Not Reversal Cursed” models (orange).

### v3

# arXiv:2509.21993v3[cs.AI] 02 Mar 2026

@@ Abstract
The reversal curse—a language model’s inability to infer an unseen fact “B is A” from a learned fact “A is B”—is widely considered a fundamental limitation. We show that this is not an inherent failure but an artifact of how models encode knowledge. Our results demonstrate that training from scratch on synthetic relational knowledge graphs leads to the emergence of a bilinear relational structure within the models’ hidden representations. This structure alleviates the reversal curse and facilitates inference of unseen reverse facts. Crucially, this bilinear geometry is foundational for consistent model editing: updates to a single fact propagate correctly to its reverse and logically dependent relations. In contrast, models lacking this representation suffer from the reversal curse and fail to generalize model edits, leading to logical inconsistencies. Our results establish that training on a relational knowledge dataset induces the emergence of bilinear internal representations, which in turn support language models in behaving in a logically consistent manner after editing. This suggests that the efficacy of language model editing depends not only on the choice of algorithm but on the underlying representational geometry of the knowledge itself.
@@ 1 Introduction
Language models have become powerful tools for knowledge-intensive tasks, yet their reasoning capabilities often fall short of human-level logical consistency (Berglund et al., 2024; Allen-Zhu & Li, 2025). A prominent example is the reversal curse: a model trained on “A is the parent of B” frequently fails to infer the reverse fact, “B is the child of A.” This failure suggests that models learn shallow, directional associations rather than robust, symmetrical relationships required for reliable inference. This limitation is particularly acute in model editing, which seeks to update factual knowledge in a trained model without costly retraining from scratch (De Cao et al., 2021; Meng et al., 2022). An ideal edit must be logically coherent; an update to the primary relation (for example, “A is the spouse of C” to “A is the spouse of D”) should naturally propagate to its logical corollaries (“D is the spouse of A” and “B is the child of D”).
However, existing architectures struggle with such logical generalization. Current model editing methods fail to propagate updates to entailed facts, requiring the explicit co-editing of both directions of a relationship to avoid the reversal curse (Thibodeau, 2022; Yao et al., 2023; Hase et al., 2024). This limitation raises a critical question: are these failures an inherent flaw in the Transformer architecture, or are they an artifact of how models represent knowledge? While prior efforts have focused on the limitations of autoregressive objectives (Zhu et al., 2024; Kitouni et al., 2024), emerging evidence indicates that these reasoning failures are tied to the specific geometry of learned representations. Language models have been shown to learn meaningful geometric structures for features like space and time (Gurnee & Tegmark, 2024; Engels et al., 2025). Furthermore, disrupting this underlying topology during editing correlates with failures in reasoning (Nishi et al., 2025). While these studies highlight that structure matters, the specific algebraic mechanism required to support symmetric and compositional reasoning remains an open question.
This work proposes that logical consistency in language models depends on the emergence of a bilinear relational structure. We investigate how relational knowledge is encoded using three representation probes: linear, translational, and bilinear probes (see Figure 1). Recent work has demonstrated that knowledge decoding from a pretrained language model’s internal representation can be approximated by either linear mapping (Hernandez et al., 2024) or translational mapping (Merullo et al., 2024). While these frameworks offer valuable insights, they cannot capture symmetric and compositional relations essential for robust logical reasoning. We instead focus on bilinear relational models such as RESCAL (Nickel et al., 2011) in knowledge graph embedding literature, which represent relations as matrices that mediate interactions between entities. Bilinear relational structures naturally accommodate inverse relations via matrix transposition and compose relations via matrix multiplication, providing a rich algebraic framework for reasoning.
@@
  • Mitigate the reversal curse via relational knowledge learning: We demonstrate that under appropriate regularization (weight decay), language models overcome the reversal curse by learning a robust bilinear relational structure, achieving near-perfect accuracy on unseen reverse relations (Figure 2, right).
  • Identify bilinear structure in hidden representations: Using three distinct probes, we find that a bilinear probe best explains hidden representations, with the highest precision emerging in the middle layers (Figure 3). Furthermore, the learned relation matrices satisfy formal algebraic tests for composition and inversion (Figure 4), validating the presence of a functional bilinear structure in language models.
  • Link representational geometry to model editing: We establish a mechanistic link between representational geometry and editing consistency. Models possessing a bilinear structure successfully propagate edits to logically related facts, whereas those lacking it fail to generalize, even when the direct edit is successful (Figure 5).
Building on these findings, our study introduces a substantial change in perspective: the key to resolving logical failures lies not solely in the model architecture or editing algorithm, but fundamentally in the geometric structure of learned knowledge representations. This result suggests the pivotal role of relational encoding in shaping the reliability and robustness of model behavior.
Figure 1: Schematics of relational embedding structures. For a subject $s$ and object $o$, the relation $r$ is modeled as: (a) linear transformation, (b) vector translation, or (c) bilinear interaction mediated by a relation-specific matrix $M_{r}$. As an example, in the fact “Son of Tracy Lance Smith is Cory Lance Smith,” $s$ is “Tracy Lance Smith,” $o$ is “Cory Lance Smith,” and $r$ is son.
@@ 2 Related work
The reversal curse—the failure to infer “B is A” from “A is B”—has been identified as a fundamental limitation of language models (Berglund et al., 2024). Prevailing hypotheses attribute this to the directional nature of the autoregressive training objective, which preferentially models $P(\text{B}|\text{A})$ but not $P(\text{A}|\text{B})$ (Allen-Zhu & Li, 2025; Zhu et al., 2024; Kitouni et al., 2024). The nature of this failure is nuanced; Lin et al. (2024) suggest that it may be an issue of knowledge retrieval rather than storage, as models that fail open-ended generation can succeed on multiple-choice questions where the answer is present in a prompt. Proposed solutions often directly target the training process. These include data augmentation via subword-level reordering in a sentence (Golovneva et al., 2024), generating reversed examples by language models (Lampinen et al., 2025), or modifying the training objective to be direction-agnostic (Kitouni et al., 2024). Although these studies offer practical mitigation, they often treat it as an unavoidable artifact of the training objective. Our work offers a different perspective by focusing on the model’s internal representations.
@@
Model editing aims to update facts in language models without costly retraining (De Cao et al., 2021). Various model editing methods have been proposed, such as ROME (Meng et al., 2022) and MEMIT (Meng et al., 2023), which show better generalization than naive fine-tuning, but all editing algorithms suffer from the reversal curse (Thibodeau, 2022; Yao et al., 2023). This points to a deeper conceptual challenge. Hase et al. (2024) provide a comprehensive critique, framing model editing as an instance of belief revision, for which no simple solution exists. They argue that logical generalization is not just a desirable feature but a core requirement for any successful editing paradigm, and they demonstrate empirically that current methods often fail to produce coherent belief updates. Recently, Nishi et al. (2025) demonstrate that editing can break the underlying topological structure of the learned knowledge. This suggests that the brittleness of model editing is tied to the preservation of internal geometry. In our work, we investigate the prerequisite internal algebraic structures that enable logical consistency after model editing.
Mathematical Structures for Relational Knowledge in Language Models.
A substantial body of research has investigated where language models store factual knowledge (Geva et al., 2021; Meng et al., 2022; Pan et al., 2025). The mechanism for retrieving specific facts is complex and distributed across multiple layers and attention heads, as revealed by studies of interventions (Hase et al., 2023) and attention mechanisms (Geva et al., 2023). In contrast, far fewer studies examine which mathematical structure models employ to resolve a relation. Recent investigations suggest that the underlying structures are unexpectedly simple: Hernandez et al. (2024) demonstrate that language models implicitly implement Linear Relational Embeddings (Paccanaro & Hinton, 2002), while Merullo et al. (2024) show that they exploit translational structures familiar from Word2Vec (Mikolov et al., 2013). However, the multi-head attention mechanism is fundamentally built on more expressive bilinear operations (Elhage et al., 2021) and individual heads can encode specific relational operations (Elhelo & Geva, 2025). Furthermore, such bilinear models have a long history of success in modeling relational data, particularly in knowledge graph embedding methods like RESCAL (Nickel et al., 2011). Despite bilinear operations being central to Transformer architecture and successful in relational learning, little work has investigated whether language models exploit bilinear structures for decoding relational knowledge. This paper presents a systematic analysis of the bilinear structures underlying relational knowledge.
@@ 3 Relational Knowledge Dataset and Language Models
In order to test our hypothesis, we create a synthetic relational knowledge dataset and train multiple language models with different hyperparameters from scratch on it.
Figure 2: (Left) Schematic of the synthetic family knowledge graph. Nodes represent entities, while directed edges represent one of eight distinct family relations used in our experimental evaluation. (Right) Test accuracy on the unseen relations (mother/father) as a function of weight decay. Each weight decay setting was trained using three different random seeds.
Synthetic Knowledge Graph and Task.
@@
Language Models.
We train decoder-only Transformers using the GPT-NeoX architecture (Andonian et al., 2023) from scratch on the synthetic dataset. Each model has 12 layers, a hidden size of 896, and 16 attention heads, totaling approximately 206M parameters. We employ this architecture in the following sections. More details of architecture and training are provided in the Appendix A.
@@ 4 Experiments
Our experiments seek to uncover the internal mechanisms that enable language models to perform logical reasoning. We first demonstrate that even when the training data contain sufficient information to infer reverse relations, models only overcome the reversal curse when guided by appropriate regularization (Experiment 1). We then pivot to the central question of our work: which mathematical structure enables this success? Through a series of probing experiments, we uncover an emergent bilinear relational structure in the successful models (Experiment 2) and verify its algebraic properties (Experiment 3). Finally, we demonstrate that this geometry plays a central role in addressing model editing challenges. We show that the bilinear structure is a key to ensuring that edits propagate in a logically consistent manner, establishing a unified explanation for both the reversal curse and editing generalization failures (Experiment 4).
4.1 Experiment 1: Training language models on relational knowledge dataset
We test whether a language model can learn to infer withheld relations (father/mother) by observing their reverse counterparts (e.g., son/daughter) and compositional examples elsewhere in the training set. We use AdamW (Loshchilov & Hutter, 2019) with a learning rate of $3\times 10^{-4}$ and sweep weight decay over $\{0,0.1,0.5,1.0,2.0,3.0,4.0,5.0,6.0\}$, training three random seeds per setting (27 models in total).
The results, shown in Figure 2 (right), reveal a striking dependency. All models achieve 100% training accuracy (see Appendix D and Figure 6), but test accuracy varies significantly based on weight decay and random seed. Without sufficient regularization, models consistently fail, showing the reversal curse with low accuracy (weight decay<~<~1.0). However, as weight decay is increased, a split in outcomes emerges: some models remain “reversal cursed,” while others break the curse and achieve near-perfect accuracy on the unseen reverse facts. This transition indicates that the reversal curse is not an inherent limitation but an artifact of an under-constrained training objective, where regularization promotes a more generalizable internal structure over simple memorization.
@@ 4.2 Experiment 2: Probing internal representation for relational structures
This probe tests for a linear relational structure (Paccanaro & Hinton, 2002) in language models. Hernandez et al. (2024) model a relation $r$ as the local affine transformation that the Transformer applies to map a subject representation $s_{l}$ from layer $l$ to the object’s pre-prediction representation $o_{L}$ in the final layer $L$. This yields the approximation $o_{L}\approx W_{r}s_{l}+b_{r}$, where $o_{L}$ is the hidden state at the position immediately preceding the object token (see Figure 1a).
@@
We estimate these parameters by averaging a small sample of $n=10$ training examples per relation. $n=10$ is chosen from a sweep over $\{10,100,500\}$ (See Appendix D.2.1).
@@
This probe, which tests for the kind of vector arithmetic investigated in hidden representations by Merullo et al. (2024), models a relation $r$ as a simple vector offset. For each relation $r$, we fit a translation vector $v_{r}$ such that $s_{l}+v_{r}\approx o_{l}$. Unlike the linear relation embedding, $o_{l}$ is taken from the same layer $l$ as $s_{l}$ (see Figure 1b) at the last token position of the object name.
@@
between any pair of entities given the relation $r$ where matrix $M_{r}\in\mathbb{R}^{d\times d}$. The target function is:
@@
Figure 3: Layer-wise probe accuracy for relational embeddings. We compare “Reversal Cursed” models (blue) against “Not Reversal Cursed” models (orange), averaged across all relations. Accuracy is reported for each layer to identify where relational knowledge emerges.
@@ Results.
For each layer $l$ and relation $r$, we trained a separate probe and evaluated its accuracy on test set entities (Figure 3). The gray dashed line at 1/3 marks the chance-level baseline. In our family graph (see Figure 2, left), for any relation $r$ there are only three candidates per family that satisfy $f^{*}_{r}(s,o)=1$. A probe that identifies the family name and the relation $r$ but fails to use the subject embedding would therefore guess uniformly among these three candidates, yielding an expected accuracy of $1/3$. For example, given the text “[Subject] Scott Wall husband [Object]”, a probe that ignores the subject can narrow the object to the three husband candidates in the “Scott Wall” family and would be correct one out of three times on average.